You own a bond with the following features: 5 years to maturity, face value of $1000, coupon rate of 2% (annual coupons) and yield to maturity of 6.3%. If you expect the yield to maturity to remain at 6.3%, what do you expect the price of the bond to be in two years? Enter the answer in dollars, rounded to the nearest cent (2 decimals).
Answer #1

Coupon =2%*Par Value =2%*1000=20
YTM =6.30%
Number of Years left = 5-2 =3
Price of bond =PV of Coupons + PV of Par Value =20*(1-(1+6.3%)^-3)/6.3%+1000/(1+6.3%)^3 =885.70